Title

Doron ZEILBERGER^{1}

\frac{U − A}{M − A} : \frac{U − B}{M − B} = \frac{M − A}{V − A} : \frac{M − B}{V − B}

1 Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA. zeilberg at math dot rutgers dot edu , http://www.math.rutgers.edu/~zeilberg . First version: xx.x, 200x. Accompanied by Maple package XXX downloadable from Zeilberger’s website. Supported in part by the NSF.